OpenStudy (anonymous):

As seen from Earth, cosmic police are traveling at 0.5c(n) while pursing bandits traveling at 0.75c(n) in order to stop the criminal, the police fire a bullet that travels at c/3(N) relative to the police. Will the bullet ever reach the bad guys?

OpenStudy (irishboy123):

as seen from earth the bullet will travel not with speed 0.5 + 1/3 = .83c but with speed \[\frac{0.5 + \frac{1}{3}}{1 + 0.5 * \frac{1}{3}} =0.71c <0.75c\]

OpenStudy (anonymous):

Thank you so much!